I need help solving these problems, I do not want the answers given to me, I want to learn step by step. Please Help.!
Show all Work
1. Solve: y^2 + 7x = 30
2. Solve: x^2 + 7 = -8x
3. Solve: x^2 – 26x + 48 = 0
4. Solve: x^2 - 4x = 32
5. Solve: w(6w + 12) = 0
6. Solve: y^2 + y = 30
7. Solve: (a – 5)(a + 3) = 0
am having big problems with 1 and 2. And for number 7 can I use FOIL?
Please answer me ASAP! Thank you very much. :)

Answer by Alan3354(69443)

1. Solve: y^2 + 7x = 30
You can't mix x & y. If you meant:
x^2 + 7x = 30
x^2 + 7x - 30 = 0
Find a pair of factors of 30 that differ by 7. One is +, one minus.
------------------------
2. Solve: x^2 + 7 = -8x
x^2 + 8x + 7 = 0
Factors of 7 are 1 & 7
-------------------------------
3. Solve: x^2 – 26x + 48 = 0
4. Solve: x^2 - 4x = 32
5. Solve: w(6w + 12) = 0
6. Solve: y^2 + y = 30
7. Solve: (a – 5)(a + 3) = 0
FOIL works